OpenStudy (anonymous):

Write the equation of the line that passes through (4,2) and is parallel to the line y=2x-1

OpenStudy (slaaibak):

Use this formula: y - y1 = m(x - x1) you know that parallel lines have the same slope, so you have m.

OpenStudy (anonymous):

you got that?

OpenStudy (anonymous):

\[y=2x-1\] looks like \[y=mx+b\] with \[m=2,b=-1\] so you know the slope of your line is \[m=2\]

OpenStudy (anonymous):

then write \[y-2=2(x-1)\] and solve for by first using the distributive property on the right and then adding 2 to both sides

OpenStudy (anonymous):

so 2x-y=?

OpenStudy (anonymous):

\[y-2=2x-2\] \[y=2x\] or if you prefer \[2x-y=0\]

OpenStudy (slaaibak):

satellite, don't you mean y - 2 = 2(x - 4) y - 2 = 2x - 8 y=2x - 6